OpenStudy (anonymous):

x + y = 5 -------------- 3x + z = 2 -------------- 4y – z = 8 x=? y=? z=?

hero (hero):

Can you re-post each equation on separate lines please?

hero (hero):

Hmmm, maybe you should post your question in this part or use the draw button. It's still not showing up right.

hero (hero):

x + y = 5 3x + z = 2 4y - z = 8

hero (hero):

Is that it?

OpenStudy (anonymous):

Yes, then find what x,y and z are.

hero (hero):

All you have to do is use substitution

hero (hero):

Isolate a variable in one of the equations

OpenStudy (dumbcow):

you can also use elimination here....get rid of "z" variable by adding 2nd and 3rd equations 3x +0y +z = 2 0x +4y -z = 8 ______________ 3x +4y = 10 Now you have system of 2 equations with 2 variables, eliminate either "x" or "y" x + y = 5 3x + 4y = 10

OpenStudy (anonymous):

A: x + y = 5 B: 3x + z = 2 C: 4y - z = 8 Cancel z using sum of B with C: 3x + z + 4y -z = 2 + 8 3x + 4y = 10 Now, isolate y: 3x + 4y = 10 4y = 10 - 3x y = (10 - 3x) / 4 Replace y in A: x + (10 - 3x) / 4 = 5 (4x + 10 - 3x) / 4 = 5 (x + 10) / 4 = 5 x + 10 = 5 * 4 x + 10 = 20 x = 20 - 10 x = 10 We get x, so now use the value of x to get y using A: x + y = 5 10 + y = 5 y = 5 - 10 y = -5 Now use the value of x to get z using B: 3x + z = 2 3*10 + z = 2 30 + z = 2 z = 2 - 30 z = -28 Check the answer using A,B and C: A: x + y = 5 10 + (-5) = 5 10 -5 = 5 5 = 5 (CORRECT) B: 3x + z = 2 3*10 + (-28) = 2 30 - 28 = 2 2 = 2 (CORRECT) C: 4y - z = 8 4*(-5) - (-28) = 8 -20 + 28 = 8 28 - 20 = 8 8 = 8 (CORRECT)